OpenStudy (anonymous):

How many solutions does the equation have? 4(2x+3)=2(3x-4) A. one solution B. infinite solutions C. no solution

OpenStudy (anonymous):

@Destinymasha

OpenStudy (stefrheart):

Have you distributed both equations?

OpenStudy (anonymous):

wouldn't it be A.? i was stuck

OpenStudy (anonymous):

It's not C.

OpenStudy (anonymous):

:O

OpenStudy (stefrheart):

Honestly just solve 4(2x+3)=2(3x-4) distribute the 4 and 2 8x+12=6x-8 subtract the 6 on both sides 2x+12=-8 subtract the 12 on both sides 2x=-20 divide the 2 x=-10 so is that one solution? infinite many? or no solutions?

OpenStudy (anonymous):

im trying to figure that out atm

OpenStudy (stefrheart):

Well did we come out with x as one number? multiple numbers? or no answer?

OpenStudy (anonymous):

i trying to find out if its A B or C

OpenStudy (stefrheart):

It should be A since there is only one x

OpenStudy (wolf1728):

4wshhs - COUNT how many numbers x equals - look at what stefrheart's answer says

OpenStudy (stefrheart):

Thanks @wolf1728

OpenStudy (wolf1728):

just thought I'd type a little something :-)

OpenStudy (anonymous):

XD

OpenStudy (stefrheart):

Hahahahaahhaha!! I was trying to put it very detailed but i love your response!

OpenStudy (anonymous):

so its B ?

OpenStudy (stefrheart):

... we just discussed this its A

OpenStudy (anonymous):

i did say A

OpenStudy (stefrheart):

ONE solution

OpenStudy (stefrheart):

you put b.. just now

OpenStudy (anonymous):

lol my bad sorry

OpenStudy (anonymous):

can u help me with another one?

OpenStudy (stefrheart):

Its okay :) Sure!

OpenStudy (wolf1728):

okay - guess I'll drop out

OpenStudy (anonymous):

thx wolf! :)

OpenStudy (anonymous):

Which expressions would complete this equation so that it has infinitely many solutions? 8 + 2(8x – 6) = Choose exactly two answers that are correct. A. 2(4x + 7) B. 9x – 10 C. 16x – 4 D. 4(4x – 1)

OpenStudy (stefrheart):

Are you sure? @wolf1728

OpenStudy (anonymous):

we might need your help :P

OpenStudy (stefrheart):

First start with simplifying the first equation

OpenStudy (stefrheart):

So what is it?

OpenStudy (anonymous):

sorry went to RR

OpenStudy (anonymous):

but im back :)

OpenStudy (anonymous):

u there @stefrheart

OpenStudy (anonymous):

i think this one is C and D honestly

OpenStudy (stefrheart):

Sorry i was studying for Chem

OpenStudy (stefrheart):

You are correct! It says chose 2 answers and C and D are correct!!! :D

OpenStudy (anonymous):

lol :D

OpenStudy (stefrheart):

Lol! Good job!

OpenStudy (anonymous):

hat grade u in ?

OpenStudy (anonymous):

what

OpenStudy (stefrheart):

Im a sophomore. You?

OpenStudy (anonymous):

9 :O

OpenStudy (anonymous):

9th

OpenStudy (stefrheart):

OH cool! Your a freshie!

OpenStudy (anonymous):

yup :D

OpenStudy (anonymous):

k stay in this room so when i need you i can message u k :DD

OpenStudy (stefrheart):

Hahahahaha!!

OpenStudy (anonymous):

lol i dont want to fail this test!

OpenStudy (stefrheart):

I know how you feel! What are you taking?

OpenStudy (anonymous):

Math atm

OpenStudy (anonymous):

How many solutions does the equation have? 3(4x-3)-7x=5x-9 A. one solution B. infinite solutions C. no solution

OpenStudy (anonymous):

k for this one im not sure on :D

OpenStudy (anonymous):

i was saying A at one point

OpenStudy (stefrheart):

Hahaha i know math but what kind?

OpenStudy (stefrheart):

Its definitely not A

OpenStudy (anonymous):

ooh algebra lol

OpenStudy (anonymous):

haha i said no to A to o

OpenStudy (stefrheart):

Oh i was confused on what you said and im in Adv. Alg. 2 Now what was your simplified versions of the 2 equations?

OpenStudy (anonymous):

it seems like they are all real numbers

OpenStudy (anonymous):

so C no?

OpenStudy (stefrheart):

If it is all real numbers it would be infinite solutions right?

OpenStudy (anonymous):

im not sure im learning from you :DD

OpenStudy (anonymous):

wait maybe it is B honestly

OpenStudy (anonymous):

I think its B

OpenStudy (stefrheart):

Lol have you not studied it before? And yes it is B infinite solutions

OpenStudy (anonymous):

lol nope just got this homework today

OpenStudy (stefrheart):

Oh did your teacher teach this?

OpenStudy (anonymous):

lol im trying to work it out before i ask you :D

OpenStudy (anonymous):

im home schooled XD this

OpenStudy (anonymous):

is a test

OpenStudy (stefrheart):

Oh that makes more sense! Now would you like for me to quickly teach you?

OpenStudy (anonymous):

Sure :D

OpenStudy (anonymous):

honestly i love to learn :D

OpenStudy (stefrheart):

One solution-there is only 1 x Infinite solutions- The equations are exactly the same and equal each other No solutions- The equations dont equal each other or one answer

OpenStudy (anonymous):

ohh

OpenStudy (stefrheart):

Exactly. Better?

OpenStudy (anonymous):

Yes better lol

OpenStudy (anonymous):

How many solutions does the equation have? 6-a=1/3(9a-18) A. one solution B. infinite solutions C. no solution

OpenStudy (anonymous):

but this one has 2 A's

OpenStudy (anonymous):

soo maaaybe

OpenStudy (stefrheart):

What do you think?

OpenStudy (anonymous):

working it out :)

OpenStudy (anonymous):

i might get it on the first try :D

OpenStudy (stefrheart):

Okay :P Thats good!!!!

OpenStudy (anonymous):

When u mean equal to eachother what do u exactly mean gotta be sure :)

OpenStudy (stefrheart):

like 5=5 or 6=2

OpenStudy (anonymous):

so is this B ?

OpenStudy (anonymous):

if not

OpenStudy (anonymous):

actuallu no its not B lol

OpenStudy (stefrheart):

Lol so its not B so it leaves A and C

OpenStudy (anonymous):

its A !

OpenStudy (anonymous):

cause for this one i cant see why it would be C lol

OpenStudy (stefrheart):

Lol you are correct it is A because a=3

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

woot! i got a 100% dude

OpenStudy (anonymous):

your AWESOME!

OpenStudy (stefrheart):

Great job! Im proud of you! I didnt do much it was mainly you! :)